We installed a new parabolic reflect mirror to focus the laser beam and adjusted the alignment of the spectrometer. In order get the signal, we changed the nozzle repetition rate to 0.5 Hz, the gas pressure decrease around 100 psi. Compared the last zeroth order spectrum, this looks very bright and sharp. The line became very weak when the nozzle was not triggered, it means that there was only 248 nm laser hit MCP.
Unfortunately, we could not find any interesting spectra when letting the spectrometer scanned in the whole range.